OpenStudy (anonymous):

Besides 0, where does the equations y=4x and 2x^2 cross? (no calculator)

OpenStudy (amistre64):

cross what?

OpenStudy (amistre64):

each other...ok right?

OpenStudy (anonymous):

yes, where the two equations cross, I will need to to plug into the bounds of green's theorem

OpenStudy (amistre64):

we want to know where y = y really and y = 4x and y = 2x^2 so..... 4x = 2x^2 when? 0 = 2x^2 -4x 0 = 2x(x-2) ; recall that any number times zero will equal zero. so either (0)(x-2) = 0 or; 2x(x-2) = 0 2x=0 when x = 0 x-2 = 0 when x = 2

OpenStudy (amistre64):

green has a theorum?

OpenStudy (amistre64):

....im pretty sure your smart enough to catch that :) 2x(0 = 0 is what I meant lol

OpenStudy (anonymous):

thanks, I hate forgetting basic factoring and algebra

OpenStudy (amistre64):

it takes practice, it like typing....you can be a pro at it and still mess it up lol

OpenStudy (anonymous):

yup thanks, I've been spoiled by a calc and my professor refuses us to ever use on in class